Assume consumption is represented by the following: C = 500 + .75Y. Also assume that planned...

  1. Assume consumption is represented by the following: C = 500 + .75Y. Also assume that planned investment (I) equals 100. Given the information, calculate the equilibrium level of income.
Answer #1

Y = C+I

Y = 500+0.75Y+100

Y = 600+0.75Y

0.25Y = 600

Equilibrium level of Income Y = 600/0.25 = 2400